5. More on debt management ratios Aa Aa E The extent of financial leverage in a firmm Debt ratios measure the proportion of total assets financed by a firms creditors Shoe Barn Inc. has a debt-to-equity ratio of 2.60, compared to the industry average of 2.08. Its competitor Heally Corp., however, has a debt-to-equity ratio of 3.90. Based on what debt-to-equity ratios imply, which of the following statements is true? O Heally Corp. has higher creditworthiness as compared to Shoe Barn Ind. O Heally Corp, has greater financial risk as compared to Shoe Barn Inc. and to the average financial risk in the industry. O Heally Corp.s creditors face lesser risk than the average financial risk in the industry. O Shoe Barn Inc.s shareholders expect magnified returns but higher risk as compared to Heally Corp. Suppose the stock pnce of Shoe Barn Inc. falls by 10%, what impact will it have on its market-to-debt ratio if nothing changes in the companys balance sheet? O The market debt ratio will increase, reflecting a decrease in the financial risk of the company. O The market debt ratio will increase, reflecting an increase in the financial risk of the company. O The market debt ratio will decrease, reflecting a decrease in the financial risk of the company. O The market debt ratio will decrease, reflecting an increase in the financial risk of the company. Data Collected S hoe Barn Inc. reported the following figures in its annual report. (Millions of dollars) Year 1
media%2F2b9%2F2b94c29c-ea48-452c-a531-53
Answer #1
  1. Heally Corp has greater financial risk as compared to Shoe Barn Inc. and to the average financial risk in the industry.

Heally Corp is much riskier than Shoe Barn Inc. as it has a higher debt to equity ratio. It shows that investors have not funded the business as much as creditors have.

  1. The market debt ratio will increase, reflecting an increase in the financial risk of the company.

The market debt ratio increases when the stock price fall by 10% and increases the risk of insolvency.

  1. Based on the information, Shoe Barn Inc. has the ability to cover its fixed financial charges 7.19 times.

It is calculated as:

EBITDA+Lease payment/Interest payment+Lease payment

=550+25/55+25= 575/80= 7.19 times.

I hope that was helpful :)